Use the Euclidean algorithm to find integers π‘₯ and 𝑦 such that 2640π‘₯ + 2110𝑦 = 10




1
Expert's answer
2021-04-28T16:35:52-0400

Let us use the Euclidean algorithm to find integers "\ud835\udc65" and "\ud835\udc66" such that "2640\ud835\udc65 + 2110\ud835\udc66 = 10". This equation is uquivalent to the equation "264\ud835\udc65 + 211\ud835\udc66 = 1".


Since

"264=211\cdot 1 +53, \ 211=53\cdot 3+52, \ 53=52\cdot 1+ 1,"

we conclude that

"1=53-52=53-(211-53\cdot 3)=-211+53\cdot 4=-211+(264-211)\cdot 4=264\cdot 4+211(-5)."


Therefore, "x=4" and "y=-5".


The general solution of the equation is "\begin{cases}
x=4-211t\\
y=-5+264t
\end{cases}"